Set solo takes 9hours 25minutes to complete a long walk. She walks (3t+2) km at 3km/h and further (t+4)km at 2km/h

1.express 9hours 25min in hours only. Leave your answer as an improper fraction

2.solve for t in the equation 9t+16 decide by 6 =113 over 12


1
Expert's answer
2021-10-08T04:00:34-0400

1) "\frac{(3t+2)km}{3km/h}+\frac{(t+4)km}{2km/h}=\frac{113hrs}{12}"


2) "12(\frac{3t+2}{3})+12(\frac{t+4}{2})=\frac{113}{12}\times 12"

"4(3t+2)+6(t+4)=113"

"12t+8+6t+24=113"

"18t=81"

"t=4.5"